391.Let a biased coin be tossed 5 times. If the probability of getting 4 heads is equal to the probability of getting 5 heads, then the probability of getting atmost two heads is:- (1)65275
- (2)5436
- (3)55181
- (4)6446
392.Four fair dice D1, D2, D3 and D4 each having six faces numbered 1,2,3,4,5 and 6 are rolled simultaneously. The probability that D4 shows a number appearing on one of D1, D2 and D3 is- (1)21691
- (2)216108
- (3)216125
- (4)216127
393.If three of the six vertices of a regular hexagon are chosen at random, then the probability that the triangle formed with these chosen vertices is equilateral is:- (1)101
- (2)51
- (3)103
- (4)203
394.Three numbers are chosen at random without replacement from {1, 2, 3, ..., 8}. The probability that their minimum is 3, given that their maximum is 6, is :- (1)83
- (2)51
- (3)41
- (4)52
395.Bag I contains 3 red, 4 black and 3 white balls and Bag II contains 2 red, 5 black and 2 white balls. One ball is transferred from Bag I to Bag II and then a ball is draw from Bag II. The ball so drawn is found to be black in colour. Then the probability, that the transferred ball is red, is:- (1)94
- (2)185
- (3)61
- (4)103
396.Let A and B be independent events such that P(A) = p,P(B) = 2p. The largest value of p, for which P (exactly one of A, B occurs)=95, is :- (1)31
- (2)92
- (3)94
- (4)125
397.A box contains 6 bottles of V1 drink, 3 bottles of V2 drink and 4 bottles of V3 drink. If three bottles are drawn at random, then the probability that the three are not of the same variety is- (1)713632
- (2)833752
- (3)858833
- (4)286261
398.A random variable 'X' has the following probability distribution : XP(X)1k−123k3k43k53k26k27k2+k Then the value of k is- (1)51
- (2)72
- (3)−2
- (4)101
399.A bag contains 6 balls. Two balls are drawn from it at random and both are found to be black. The probability that the bag contains at least 5 black balls is- (1)75
- (2)72
- (3)73
- (4)47
400.It is given that the discrete random variable X∼B(n,p) and P(X=2)=P(X=3). Then the mean of X is- (1)2−p
- (2)3−p
- (3)p−2
- (4)p−3
401.If A and B are two events such that P(A∪B)=65, P(Aˉ)=41 and P(B)=31, then A and B are- (1)mutually exclusive
- (2)independent events
- (3)exhaustive events
- (4)exhaustive and independent events
402.Let Ω be the sample space and A⊆Ω be an event. Given below are two statements: (S1) : If P(A)=0, then A=ϕ (S2) : If P(A)=1, then A=Ω Then- (1)Only (S1) is true
- (2)Only (S2) is true
- (3)both (S1) and (S2) are true
- (4)Both (S1) and (S2) are false
403.For two events A and B, a true statement among the following is- (1)P(A′∪B′)=1−P(A)P(AB)
- (2)P(A′∪B′)=1−P(A∪B)
- (3)P(A′∪B′)=P(A∪B)
- (4)P(A′∪B′)=P(A′)+P(B′)
404.The probability distribution of a random variable X is given below: X: 4k, 730k, 732k, 734k, 736k, 738k, 740k, 6k P(X): 152, 151, 152, 51, 151, 152, 51, 151 If E(X)=15263, then P(X<20) is equal to :- (1)1511
- (2)53
- (3)1514
- (4)158
405.Let A,B,C be three pairwise independent events of a random experiment. If P(Bˉ∪Cˉ)=21, P(A)>0, P(B)=b and P(C)=c, then P((Bˉ∩Cˉ)/A)=- (1)1+b−c
- (2)2+b−c
- (3)23−b−c
- (4)2−b−c
406.For the random variable X with probability distribution given by the table X=x: 0, 1, 2, 3 P(X=x): K, K+71, 2K, 52 The mean of X is- (1)3531
- (2)3557
- (3)3563
- (4)3567
407.If X has a binomial distribution with parameters n=6, p and P(X=2):P(X=3)=12:5, then p=- (1)125
- (2)21
- (3)2116
- (4)215
408.In a city, 10 accidents take place in a span of 50 days. Assuming that the number of accidents follow the Poisson distribution the probability that three or more accidents occur in a day, is- (1)∑k=3∞k!e−λλk, λ=0.2
- (2)∑k=3∞k!eλλk,λ=0.2
- (3)1−∑k=03k!e−λλk,λ=0.2
- (4)∑k=03k!e−λλk,λ=0.2
409.If A and B are two independent events such that P(Aˉ)=0.75, P(A∪B)=0.65 and P(B)=x, then find the value of x.- (1)145
- (2)149
- (3)158
- (4)157
410.Let A and B be two events with P(A)=71, P(BA)=52 and P(B)=72. Then, the value P(AB) is- (1)51
- (2)495
- (3)54
- (4)53
411.If A and B are two events such that P(A∩B)=0.1, and P(A∣B) and P(B∣A) are the roots of the equation 12x2−7x+1=0, then the value of P(A∩B)P(A∪B) is:- (1)34
- (2)47
- (3)35
- (4)49
412.If E and F are events such that P(F′)=0.7 and P(E∩F)=0.2, then P(E/F) is- (1)2/3
- (2)1/3
- (3)3/4
- (4)1/4
413.If S be the sample space of a random experiment ξ and P be a probability function defined on the power set P(S) of S, then which one of the following is not satisfied by P? (i) P(ϕ)=0 (ii) If Ec is the complementary event of E, then P(Ec)=1−P(E) (iii) 0≤P(E)≤1,∀E⊆S (iv) If E1⊆E2P(E2)≤P(E1)- (1)(iii)
- (2)(iv)
- (3)(ii)
- (4)(i)
414.In a group of 400 people, 160 are smokers and non-vegetarian; 100 are smokers and vegetarian and the remaining 140 are non-smokers and vegetarian. Their chances of getting a particular chest disorder are 35%, 20% and 10% respectively. A person is chosen from the group at random and is found to be suffering from the chest disorder. The probability that the selected person is a smoker and non-vegetarian is:- (1)457
- (2)458
- (3)4514
- (4)4528
415.A shopkeeper buys a particular type of electric bulbs from three manufactures M1,M2 and M3. He buys 25% of his requirement from M1,45% from M2, and 30% from M3. Based on past experience he found that 2% of type M3 bulbs are defective, where as only 1% of type M1 and type M2 are detective. If a bulb chosen by him at random is defective, then the probability that it was of type M3 is- (1)135
- (2)136
- (3)137
- (4)138
416.Cards are drawn one after the other without replacement from a well shuffled pack of cards until and ace card appears. If the probability that exactly 5 cards are drawn before the first ace card appears is 494(p4⋅p5⋅p6p1⋅p2⋅p3),(pi) is prime, i=1,2,3,4,5,6) then (max(pi)−min(pi))=- (1)12
- (2)18
- (3)20
- (4)22
417.A company has two plants A and B to manufacture motorcycles. 60% motorcycles are manufactured at plant A and the remaining are manufactured at plant B.80% of the motorcycles manufactured at plant A are rated of the standard quality, while 90% of the motorcycles manufactured at plant B are rated of the standard quality. A motorcycle picked up randomly from the total production is found to be of the standard quality. If p is the probability that it was manufactured at plant B, then 126p is- (1)54
- (2)66
- (3)64
- (4)56
418.If the probability that the random variable X takes values x is given by P(X=x)=k(x+1)3−x, x=0,1,2,3,…, where k is a constant, then P(X≥2) is equal to- (1)277
- (2)187
- (3)1811
- (4)2720
419.The minimum number of times a fair coin needs to be tossed, so that the probability of getting at least two heads is at least 0.96, is 420.Of the three independent events E1, E2 and E3, the probability that only E1 occurs is α,only E2 occurs is β and only E3 occurs is γ. Let the probability p that none of events E1, E2 or E3 occurs satisfy the equations (α – 2β) p = αβ and (β – 3γ) p = 2βγ . All the given probabilities are assumed to lie in the interval (0, 1). Then ProbabilityofoccurrenceofE3ProbabilityofoccurrenceofE1=
|